OpenStudy (anonymous):

The equation of the line that goes through the points (5,9) and (6,3) can be written in the form y=mx+b where m is? B is?

OpenStudy (australopithecus):

m = slope b = y intercept use the equation m = (y2 - y1)/(x2-x1)

OpenStudy (australopithecus):

where a point is (x,y) once you have m insert one of the points into the function: so y = mx + b becomes 3 = m(6) + b and solve for b

OpenStudy (australopithecus):

(x1,y1), (x2, y2) (5, 9), (6 , 3) Insert these values into the formula I provided for m

OpenStudy (anonymous):

slope is just rise over run or the change in y values over the change in x values. which means that m which is the slope is equal to (y2-y1)/(x2-x1). once you find the slope you can construct a y=mx+b formula using one of the given point to plug in your x and y values. with this you can solve for b. once you have b and m you just rewrite the formula again plugging in only the m and b values and that will be your formula for the line

OpenStudy (anonymous):

Thank u 4 that expalnation! I really apprici8 it!